OpenStudy (anonymous):

If = 27° and x = 9 cm, what is the value of y to the nearest tenth of a centimeter? alpha is equal to 27 degrees by the way

OpenStudy (anonymous):

What equals 27 degrees? I think you forgot to put it in

OpenStudy (anonymous):

Do you have any other info? Is there a circle involved or something?

OpenStudy (anonymous):

so its a triangle with a right angle on the bottom right, alpha is at the very top, theta is at the bottom left, x is on the bottom outside horizontal line and z is outside on the left side (hypotenuse) and y is outside on the right

OpenStudy (anonymous):

ok, thanks :) one moment

OpenStudy (anonymous):

could u write ur steps down as well

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok. So The first thing you want to do when solving a triangle is to identify how many sides and angles you have and in what pattern. It is AAS (angle-angle-side in this case. Each side always coordinates with the angle across from it. So x coordinates with 27 degrees (X), etc. Now, use the Law of Sines to find the side y: y/ sin Y = x/ sin X y/ sin Y = 9/ sin 27 We need to find the angle Y, which is directly across from side y. This part is easy: add 27 + 90 (the angles you have) = 117. Then subtract 117 from 180 = 63. y/ sin 63 = 9/ sin 27 Use a calculator to find the sines of 63 and 27: y/ 0.8910 = 9/ 0.4540 Now multiply both sides by 0.8910 to isolate x: y = (9/0.454) * (0.891/1) y = 8.019/0.454 y = 17.6630 = 17.7

OpenStudy (anonymous):

a good reference: http://www.mathsisfun.com/algebra/trig-solving-triangles.html

OpenStudy (anonymous):